Marked price of a commodity is 35% above the cost price. If he gives a discount of 15%, how much he gains on the deal.

Grade:

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Hemant

let actual price is 100

marked price=135

discount=135*0.15=20.25

selling price=135-20.25=114.75

profit is 14.75%

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Hemant

cost price=100

marked=135

discount=135*0.15=20.25

selling price=135-20.25=114.75

net gain 14.75%

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..